two types of fertilizer are being used, one costs 14 dollars per pound and the other is 18 dollars per pound. if the mixture endes with a total of 200 pounds and costs 15 dollars per pound how much of each type is used?
Answer by Maths68(1474)

Let
Type 1 = x pound
Type 2 = 200-x pound
Mixture = 200 pound
=============================================================================
Type 1 cost per pound = 14$
Type 2 cost per pound = $18
Mixture cost per pound = $15
==============================================================================
(Cost of Type 1 * Type 1)+(Cost of Type 2 * Type 2)=(Cost of Mixture * Mixture)
(14*x)+(18*(200-x))=15*200
14x+3600-18x=3000
14x-18x=3000-3600
-4x=-600
-4x/-4=-600/-4
x=150
=========================================================================
Type 1 = x = 150 pound
Type 2 = 200-x=200-150= 50 pound
Mixture = 200 pound
=========================================================================
Check
(150*14)+(50*18)=200*15
2100+900=3000
3000=3000